# Differential Privacy of Cross-Attention with Provable Guarantee

### v2

@@ I.4 Proof of Main Result
Theorem I.12 (Softmax cross-attention, formal version of Theorem 3.1).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Let $\alpha\in(0,1)$ be the relative error parameter and $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma C.7). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition I.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeSoftmaxAdaptive (Algorithm 1) that uses $O(lnrd)$ spaces to make cross-attention DP and supports the following operations:
  • We initialize $d$ data structures using Init$(K,n,V_{*,k},\epsilon\in(0,1),\delta\in(0,1),\delta^{\prime}\in(0,1),c\in(0,0.1),\epsilon_{s}\in(0,0.1),p_{f}\in(0,0.01))$ (Algorithm 1), for $k\in[d]$. It takes $O(lnr)$ time to initialize one data structure.
  • At query time, for user input $Q$, we process one token at a time by passing the $i$-th row of $Q$, denoted $Q_{i}\in\mathbb{R}^{d}$, to DistanceQuery$(Q_{i},\alpha\in(0,1))$ (Algorithm 1) for each $i\in[m]$. It takes $O(\alpha^{-1}lr\log^{2}n)$ time to output an entry $z$ in $\mathrm{Attn}(Q,K,V)$ such that
@@
    • the process of output $z$ has relative error $n^{-1}(\alpha+\epsilon_{s})$,
    • the process of output $z$ has additive error $O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n)$,
@@
From Section 3, we know that we can ensure matrix $AV$ in cross-attention computation satisfies DP. Next, from Theorem 4.4, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query. For $D$, we compute the exact true value. By the post-processing property of DP (Fact B.5), our cross-attention process is DP.
Let $(AV)_{i,k}$ be the true value and $\widetilde{(AV)}_{i,k}$ be the noisy value. Then, from Theorem 4.4, we have
$|(AV)_{i,k}-\widetilde{(AV)}_{i,k}|\leq(\alpha+\epsilon_{s})\cdot(AV)_{i,k}+O(\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n).$ (11)
@@
Finally, we can show the error of one entry is bounded by
@@
$\leq$ $~{}n^{-1}(\alpha+\epsilon_{s})\cdot(AV)_{i,k}$
$~{}+O(n^{-1}\epsilon^{-1}\alpha^{-1/2}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}(n))$
where the first step follows from definition, the second step follows from simple algebra, and the last step follows from Eq.(11) and (12). ∎

### v3

@@ G.4 Proof of Main Result
Theorem G.11 (Softmax cross-attention, formal version of Theorem 4.1).
Let $Q,K,V,\mathrm{Attn}$ be defined in Definition 1.1. Assume the input context length $n$ is large enough. Let $p_{f}$ be the probability of failure parameter. Let $r,s,\epsilon_{s}$ be parameters of polynomial kernel methods (Lemma G.6). Let $\Gamma_{R,s}:=\max_{j\in[s]}\frac{R^{j}}{\sqrt{j!}}$ (Definition G.3). Let $l=O(r\log(dR/(\epsilon_{s}p_{f})))$. There is a data structure DPTreeCrossAttention (Algorithm 1) that uses $O(lnrd)$ spaces to ensure cross-attention DP and supports the following operations:
[…]
@@
    • the process of output $z$ has relative error $2\epsilon_{s}/(1-\epsilon_{s})$,
    • the process of output $z$ has additive error $O((1-\epsilon_{s})^{-1}n^{-1}\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n),$
@@
We first prove the privacy and then prove error for each coordinate of the output $O$ of Algorithm 1.
Proof of Privacy:
From Theorem G.10, $\mathcal{D}_{k}$.DistanceQuery for $k\in\{0,1,\dots,d\}$ in Algorithm 1 answer $(\epsilon/2,\delta/2+\delta^{\prime}/2)$-DP queries that are robust to adaptive queries. By Fact A.8, the procedure for calculating each coordinate of vector $O$ is $(\epsilon,\delta+\delta^{\prime})$-DP in Line 15 of Algorithm 1.
Proof of Error:
We prove the error bound of the cross-attention module. We omit the constant factor of 2 used for the privacy guarantee in Algorithm 1, incorporating it into the big-$O$ notation for the error analysis. Let $AV$ be the true value and $\widetilde{AV}$ be the noisy value. Let $D$ be the true value and $\widetilde{D}$ be the noisy value. First, we use triangular inequality to decompose the error:
$~|(D^{-1}AV)_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
$\leq$ $~|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|+|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$ (11)
We now prove for each term.
Part 1: Error bound for AVAV
From Section 4, we know that we can ensure matrix $AV$ in cross-attention computation satisfies DP. Next, from Theorem 5.3, for $i\in[m],j\in[n],k\in[d]$, we have $(AV)_{i,k}$ is $(\epsilon,\delta+\delta^{\prime})$-DP and also robust to adaptive query.
Let $\zeta:=\epsilon^{-1}l\Gamma_{R,s}^{2}R_{w}r\sqrt{\log(l/\delta^{\prime})}\cdot\log^{3/2}n$ denote the additive error. Then, from Theorem G.10, we have
$|(AV)_{i,k}-\widetilde{(AV)}_{i,k}|\leq$ $~|\epsilon_{s}\cdot(AV)_{i,k}|+O(\zeta)$ (12)
@@
Finally, we can show the error of first term in Eq. (G.4) is bounded by
@@
$\leq$ $~|\epsilon_{s}\cdot D_{i,i}^{-1}(AV)_{i,k}|+O(n^{-1}\zeta)$
where the first step follows from definition, the second step follows from simple algebra, and the last step follows from Eq. (12) and (13).
Part 2: Error bound for $D$
[…]
Recall that
$D_{i,i}=\sum_{i=1}^{n}\exp(\langle Q_{i},K_{j}\rangle/d)).$
This can be viewed as the weighted Softmax problem but with weight ${\bf 1}_{n}$. To be more clear, let us recall that $R_{w}$ is the upper bound of the entries in $V$, and define $R_{w}^{\prime}$ as the upper bound of the entries in ${\bf 1}_{n}$. Observe that we can reuse previous results in Theorem G.10 with adjustment only on the value of $R_{w}^{\prime}$ (which is 1) in $\mathcal{D}_{0}$.
We wish to bound
$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|=$ $~\frac{|D_{i,i}-\widetilde{D}_{i,i}|}{D_{i,i}\cdot\widetilde{D}_{i,i}}.$
For the term $|D_{i,i}-\widetilde{D}_{i,i}|$, similar to Eq. (12), from Theorem G.10, we have
$|D_{i,i}-\widetilde{D}_{i,i}|\leq|\epsilon_{s}\cdot D_{i,i}|+O(\zeta),$ (14)
where we assume $R_{w}\geq 1=R_{w}^{\prime}$ and loose the $R_{w}^{\prime}$ in additive error parameter in $\mathcal{D}_{0}$ from 1 to $R_{w}$.
Now we need the lower bound of $\widetilde{D}_{i,i}$. From Eq. (14), we have
$\widetilde{D}_{i,i}\geq$ $~D_{i,i}-(|\epsilon_{s}\cdot D_{i,i}|+O(\zeta))\geq|(1-\epsilon_{s})\cdot D_{i,i}|-O(\zeta).$
Then, we have
$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|=$ $~D_{i,i}^{-1}\frac{|D_{i,i}-\widetilde{D}_{i,i}|}{\widetilde{D}_{i,i}}\leq D_{i,i}^{-1}\frac{|\epsilon_{s}\cdot D_{i,i}|+O(\zeta)}{|(1-\epsilon_{s})\cdot D_{i,i}|-O(\zeta)}$
We assume $n$ is large enough and thus ignore other small factors. Observe that $O(\zeta)=O(\log^{3/2}n)$, and $D_{i,i}\geq n=O(n)$ from Part 1. Thus, $O(\zeta)$ is a small order term compared to $D_{i,i}$. As a consequence, we get
$|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|\leq D_{i,i}^{-1}\frac{|\epsilon_{s}\cdot D_{i,i}|}{|(1-\epsilon_{s})\cdot D_{i,i}|}=D_{i,i}^{-1}\frac{\epsilon_{s}}{(1-\epsilon_{s})},$ (15)
since $\epsilon_{s}\in(0,0.1)$.
From Eq. (12), we have
$|\widetilde{(AV)}_{i,k}|\leq$ $~(1+\epsilon_{s})\cdot|(AV)_{i,k}|+O(\zeta)$
We consider the second term in Eq.(G.4). Then,
$~|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
$=$ $~|D^{-1}_{i,i}-\widetilde{D}^{-1}_{i,i}|\cdot|(\widetilde{AV})_{i,k}|$
$\leq$ $~D_{i,i}^{-1}\frac{\epsilon_{s}}{(1-\epsilon_{s})}((1+\epsilon_{s})\cdot|(AV)_{i,k}|+O(\zeta))$
$=$ $~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot D_{i,i}^{-1}|(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}D_{i,i}^{-1}\zeta)$
$\leq$ $~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}n^{-1}\zeta)$
where the first step follows from simple algebra, the second step follows from the previous derived upper bounds, the third step follows from simple algebra, and the last step follows from Eq.(13).
Part 3: Final error bound
Combining results from Part 1 and 2, the final error bound is
$~|(D^{-1}AV)_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
$\leq$ $~|(D^{-1}AV)_{i,k}-(D^{-1}\widetilde{AV})_{i,k}|+|(D^{-1}\widetilde{AV})_{i,k}-(\widetilde{D}^{-1}\widetilde{AV})_{i,k}|$
$=$ $~\epsilon_{s}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(n^{-1}\zeta)+~\epsilon_{s}\frac{(1+\epsilon_{s})}{(1-\epsilon_{s})}\cdot|D_{i,i}^{-1}(AV)_{i,k}|+O(\frac{\epsilon_{s}}{(1-\epsilon_{s})}n^{-1}\zeta)$
$=$ $~\frac{2\epsilon_{s}}{(1-\epsilon_{s})}\cdot|(D^{-1}AV)_{i,k}|+O((1-\epsilon_{s})^{-1}n^{-1}\zeta)$
Therefore, we prove the error bound.